OpenStudy (anonymous):

Simplify 11(sqrt)112

OpenStudy (anonymous):

\[11\sqrt{112}\]

OpenStudy (anonymous):

you need to decompose 112 into factors. 112 = 4* 4* 7, right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so now you have 11*sqrt(4*4*7). How could you simplify what's in the square root?

jhonyy9 (jhonyy9):

11sqrt(2*8*7)=11sqrt(16*7)=11*4sqrt7 =44sqrt7

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

\[\sqrt{16*7} = \sqrt{16}*\sqrt{7}\] Is it more visible now?

OpenStudy (anonymous):

\[A. -44\sqrt{7}\] \[B. -176\sqrt{7}\] \[C. -26\sqrt{7}\] \[D. 4\sqrt{7}\]

OpenStudy (anonymous):

the sqrt of 16 is 4, so it is D?

OpenStudy (anonymous):

I forgot to put the 11 in there, so it's actually A) : 11 * (4sqrt(7))

OpenStudy (anonymous):

But, how is it -44? shouldint it be positive?

OpenStudy (anonymous):

sqrt(16) is either 4 or -4, because 4^2 = 16 and (-4)^2 = 16. Usually, we always deal in positives, but the sqrt of something can be either sign. Important, what's INSIDE the sqrt root sould always be positive tho.

OpenStudy (anonymous):

Ohhh ok. thanks. :)